find the resultant of two forces of 4N and 5N is. 1. they are inclined perpendicular to each other. 2. they are inclined at 350 to each other. 3. they are inclined at 1450 to each other


1
Expert's answer
2021-04-05T11:09:10-0400

1)


"R=\sqrt{4^2+5^2+2(4)(5)\cos{90}}=6.4\ N"

2)


"R=\sqrt{4^2+5^2+2(4)(5)\cos{35}}=8.6\ N"

3)


"R=\sqrt{4^2+5^2+2(4)(5)\cos{145}}=2.9\ N"